# Estimating Dense-Packed Zone Height in Liquid-Liquid Separation: A Physics-Informed Neural Network Approach

### v1

@@ 5 Phase height estimation
We now apply the models to estimate the phase heights in the gravity settler. In Section˜5.1 we present the state estimation framework that is inspired by the EKF to track and estimate the settler heights. Next, we describe the simple neural network to predict the DPZ height at separator outlet in Section˜5.2. Finally, in Section˜5.3, we discuss the estimation results, comparing PINN and VNN.
@@ 5.1 State estimation framework
where $\mathbf{F}_{k}=\frac{\partial\bm{f}_{\bm{\theta}}}{\partial\mathbf{\mathbf{x}}}\big|_{\hat{\mathbf{x}}_{k-1|k-1}}$ is the Jacobian of the state transition model $\bm{f}_{\bm{\theta}}(\cdot)$ with respect to the estimated states, $\mathbf{P}_{k}$ is the state covariance matrix, and $\mathbf{W}_{k}$ is the process noise covariance at time step $k$. Here, we initialize $\mathbf{P}_{0}$ as a diagonal matrix with values 0.0001 and 0.0001, corresponding to $\mathbf{x}=[h_{\textrm{HP}},h_{\textrm{DP}}]^{\top}$. The process noise covariance $\mathbf{W}_{k}$ is computed adaptively at each time step. Such adaptive sampling of the covariance matrix $\mathbf{W}_{k}$, sometimes referred to as Adaptive Kalman Filtering (mohamed1999adaptive), is a well-established technique that has been shown to improve both the stability and performance of the Kalman Filter (mohamed1999adaptive; zhang2020identification). We leverage the model ensemble of 40 independently trained PINNs to calculate the process noise covariance $\mathbf{W}_{k}$ in the following way: (i) First, we take the ensemble mean of the last estimated state $\hat{\mathbf{x}}_{k-1|k-1}$ and feed it as the initial state to the transition model $\bm{f}_{\bm{\theta}}(\cdot)$ of each ensemble member individually. (ii) We then obtain the next state estimate $\hat{\mathbf{x}}_{k|k-1}$ from each ensemble member. (iii) We calculate the sample covariance of these 40 predictions to obtain the process noise covariance $\mathbf{W}_{k}$ at each time step $k$.
@@
where $\mathbf{H}_{k}=\frac{\partial\bm{h}_{\bm{\theta}}}{\partial\mathbf{x}}\big|_{\hat{\mathbf{x}}_{k|k-1}}$ is the Jacobian of the measurement model $\bm{h}_{\bm{\theta}}(\cdot)$ w.r.t to the estimated states. The notation $\hat{\mathbf{x}}_{k|k}$ represents the updated state at time step $k$ after incorporating the measurement at that step. The measurement noise covariance matrix $\mathbf{R}$ is diagonal and constant, with entries derived from the flow sensor specifications. Since both flow measurements come from the same sensor, a single $\mathbf{R}_{k}$ value is used for both states. The sensor specifies a maximum error of $e_{\mathrm{max}}=\pm 0.15\%$. Assuming a Gaussian error distribution, the standard deviation is approximated using the empirical $3\sigma$ heuristic as $\sigma=e_{\mathrm{max}}/3=5\times 10^{-4}$ (wackerly2008mathematical). The resulting covariance is $\mathbf{R}_{\mathrm{sensor}}=\sigma^{2}=2.5\times 10^{-7}$. Note that we use the Joseph form (bucy2005filtering) for the covariance update in Eq.˜14, for improved stability. Preliminary analysis of the validation trajectory showed that the average ratio of $\frac{\mathbf{W}_{k}}{\mathbf{R}_{k}}$ was high, causing the filter to overtrust measurement updates and yield noisy state estimates. To address this, $\mathbf{W}_{k}$ was decreased by a factor of 100. After each update, the PINN initial state is reset to $\hat{\mathbf{x}}_{k|k}$ to ensure that the state transition model starts from the most recent state estimate. We apply the filter for each model in the ensemble separately, and we only use the ensemble mean of state estimates for the process noise covariance calculations.
@@ 5.3 Results
This section presents the results of the state estimation analysis. First, an initial guess for the phase heights is obtained by uniformly sampling 100 height vectors within their bounds and using those as initial state inputs to PINN (and VNN) to predict the outlet volume flows at $\tau=0$. These predicted outflows are compared against the experimentally measured outlet flows at $\tau=0$, and the candidate yielding the best agreement is selected as the initial state estimate. We then feed these initial states into the filter and proceed to track and estimate the phase heights for each PINN (and VNN) model individually. The individual trajectories, along with the ensemble means, are shown in Figs. 11 and 12.
In Fig.˜11, we report the results for the interpolation test trajectory. For both PINN and VNN, the estimated initial DPZ height $h_{\textrm{DP}}(\tau_{0})$ deviates from the experimental trajectory, indicating weak reconstruction of the internal states from outlet flow measurements alone. Nevertheless, in both cases the ensemble mean converges to the experimental trajectory at around $\tau=$100\text{\,}\mathrm{s}\text{/}$\$. For the heavy phase height, the PINN consistently outperforms the VNN. Despite the significant overshoot observed in the forward simulation results shown in Fig.˜9(b), the VNN–filter combination achieves comparable accuracy to the PINN ensemble in predicting DPZ height for the interpolation trajectory. This indicates that the filter update step effectively compensates for the VNN’s error accumulation by incorporating outflow measurements. However, compared to the PINN, the individual estimates of the VNN ensemble exhibit lesser consistency.
For the extrapolation test trajectory in Fig.˜12, the PINN ensemble performs better than the VNN ensemble in case of the DPZ height. In particular, after the maximum control action (an operating value unseen in the training trajectory) is applied at around process time $\tau=$500\text{\,}\mathrm{s}\text{/}$\$, the VNN mean estimate overshoots the experimental DPZ height, whereas the PINN continues to track the DPZ height with good accuracy. Again, the individual estimates from the PINN ensemble are also more consistent than the ones from the VNN. For the heavy phase height, the VNN and PINN ensembles demonstrate similar accuracy, with the PINN underestimating the heavy phase height around process time $\tau=$500\text{\,}\mathrm{s}\text{/}$\$, while the VNN shows increasing deviation from the experimental trajectory thereafter.
@@
Figure 11: State estimation results for the interpolation test trajectory. Estimations from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line.
@@
Figure 12: State estimation results for the extrapolation test trajectory. Estimations from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line.
After estimating the average DPZ height along the separator using the filter, we apply the simple NN to predict the DPZ height at the separator end (Fig.˜13). Both the PINN and VNN ensembles demonstrate good accuracy. Deviations occur primarily when the average DPZ height itself deviates, indicating that the NN performs well, but errors propagate where the underlying ensemble predictions are inaccurate (cf. Figs.˜11 and 12). This effect is emphasized in the extrapolation trajectory (cf. Fig.˜12), where larger deviations in the VNN ensemble lead to correspondingly larger errors in the predicted DPZ height at the separator end.
@@
Figure 13: Prediction of the DPZ height at the separator end. Experimental results are obtained from camera QIR04 at detection window $h_{4,3}$. Estimates of the average DPZ from both PINN and VNN are fed into the NN that predicts the DPZ height at separator end.
We therefore conclude that the two-stage trained PINN, when combined with the filter, enables effective real-time estimation of the internal separator states based only on outlet flow measurements. As the outlet flow rate measurements are inexpensive to obtain, our PINN constitutes a practical soft sensor for the difficult-to-measure phase heights. Moreover, we demonstrate that predicting the DPZ height at the separator end (the maximum value) from the average DPZ height is possible using a simple NN.
@@ 6 Conclusion and Outlook
We investigated the estimation of phase heights in a pilot-scale liquid–liquid separator operating with 1-octanol and deionized water at 30 °C under varying flow conditions. Direct measurement of phase heights is challenging, as the interior of the separator is not readily accessible for reliable observation of the phase interface. Although we employ a YOLOv8-based phase-boundary detection algorithm in our experimental setup, the resulting measurements still contain noise due to the difficulty of distinguishing the interface and the sensitivity of the algorithm to illumination conditions. Moreover, the available mechanistic model relies on multiple simplifying assumptions, resulting in limited predictive accuracy. To address these challenges, we developed a PINN-based dynamic model and proposed a two-stage training strategy: pretraining on approximate physics, i.e., synthetic data from the low-fidelity model, followed by fine-tuning with experimental data. We further used ensemble learning to mitigate the influence of outlier models. Finally, we combined the trained PINN with an EKF-inspired filter to enable estimation of the aqueous phase and DPZ heights. The filter prevents error accumulation from chaining of PINN predictions by fusing model predictions with real-time flow measurements to update the estimated phase-height trajectories.
@@
The two-stage PINN ensemble outperformed both the low-fidelity mechanistic model and a single-stage PINN ensemble trained only on experimental data, showing that integrating mechanistic model knowledge with experimental measurements yields better results than either approach alone. For further comparison, we also pretrained and fine-tuned an ensemble of purely data-driven vanilla neural networks (VNNs). Both the PINN and VNN ensembles were embedded in a state estimation framework to enable phase height estimation from flow rate measurements only. The PINN ensemble achieved higher accuracy in tracking phase heights. Furthermore, a simple neural network can predict the outlet DPZ height (typically the maximum value and most critical for flooding) from average DPZ height estimates. The results demonstrate that the PINN-based state estimation can serve as a reliable soft sensor, enabling the real-time monitoring of this separation process from readily available flow rate measurements.
@@
A few limitations of the present study should be noted. First, the model assumes a band-shaped DPZ, which becomes inaccurate at higher flow rates where the DPZ develops a distinct wedge-shaped profile. Although a mapping between the average and maximum DPZ height was learned by a simple neural network, the PINN model does not fully represent the actual separator dynamics. Second, the experimental data contain substantial gaps caused by missing phase-boundary detections, particularly for QIR03, which limits the training data quality. Third, the experimental dataset includes no trajectories with conditions close to flooding. While such conditions are difficult to obtain experimentally for safety reasons, their absence restricts the ability of the model to detect or classify flooding behavior. Lastly, we acknowledge that filtering the immeasurable internal flows in the proposed state estimation framework could improve phase height estimation accuracy, however this is not possible with the current PINN architecture.
@@ Acknowledgements
During the preparation of this work, generative AI tools (GPT-5) were used to assist with grammar, spelling, and style improvements in the writing. Generative AI coding tools (Claude 4) were used for post-processing tasks, specifically for plotting and visualization of the results. In all cases, the authors reviewed and edited the outputs, and they take full responsibility for the integrity and accuracy of the publication.

### v2

@@ 5 Phase height estimation
We now apply the models to estimate the phase heights in the gravity settler. It is important to distinguish between training and deployment: While phase height measurements are used during training, they will not be available during deployment. In Section˜5.1 we present the state estimation framework that is inspired by the EKF to track and estimate the settler heights. Next, we describe the simple neural network to predict the DPZ height at separator outlet in Section˜5.2. Finally, in Section˜5.3, we discuss the estimation results, comparing PINN and VNN.
@@ 5.1 State estimation framework
where $\mathbf{F}_{k}=\frac{\partial\bm{f}_{\bm{\theta}}}{\partial\mathbf{\mathbf{x}}}\big|_{\hat{\mathbf{x}}_{k-1|k-1}}$ is the Jacobian of the state transition model $\bm{f}_{\bm{\theta}}(\cdot)$ with respect to the estimated states, $\mathbf{P}_{k}$ is the state covariance matrix, and $\mathbf{W}_{k}$ is the process noise covariance at time step $k$. Here, we initialize $\mathbf{P}_{0}$ as a diagonal matrix with values 0.0001 and 0.0001, corresponding to $\mathbf{x}=[h_{\textrm{HP}},h_{\textrm{DP}}]^{\top}$.
The process noise covariance $\mathbf{W}_{k}$ is computed adaptively at each time step. Such adaptive sampling of the covariance matrix $\mathbf{W}_{k}$, sometimes referred to as Adaptive Kalman Filtering (mohamed1999adaptive), is a well-established technique that has been shown to improve both the stability and performance of the Kalman Filter (mohamed1999adaptive; zhang2020identification). We leverage the model ensemble of 40 independently trained PINNs to calculate the process noise covariance $\mathbf{W}_{k}$ in the following way: (i) First, we take the ensemble mean of the last estimated state $\hat{\mathbf{x}}_{k-1|k-1}$ and feed it as the initial state to the transition model $\bm{f}_{\bm{\theta}}(\cdot)$ of each ensemble member individually. (ii) We then obtain the next state estimate $\hat{\mathbf{x}}_{k|k-1}$ from each ensemble member. (iii) We calculate the sample covariance of these 40 predictions to obtain the process noise covariance $\mathbf{W}_{k}$ at each time step $k$. Here, the ensemble spread is interpreted as a sample-based approximation of the prediction uncertainty of the PINN model. The sample covariance of the ensemble predictions is used as a proxy for the process noise covariance $\mathbf{W}_{k}$, thereby incorporating model (epistemic) uncertainty into the filtering procedure. This allows the filter to account for increased uncertainty in regions where the model predictions diverge.
@@
where $\mathbf{H}_{k}=\frac{\partial\bm{h}_{\bm{\theta}}}{\partial\mathbf{x}}\big|_{\hat{\mathbf{x}}_{k|k-1}}$ is the Jacobian of the measurement model $\bm{h}_{\bm{\theta}}(\cdot)$ w.r.t to the estimated states. The notation $\hat{\mathbf{x}}_{k|k}$ represents the updated state at time step $k$ after incorporating the measurement at that step. The measurement noise covariance matrix $\mathbf{R}$ is diagonal and constant, with entries derived from the flow sensor specifications. Since both flow measurements come from the same sensor, a single $\mathbf{R}_{k}$ value is used for both states. The sensor specifies a maximum error of $e_{\mathrm{max}}=\pm 0.15\%$. Assuming a Gaussian error distribution, the standard deviation is approximated using the empirical $3\sigma$ heuristic as $\sigma=e_{\mathrm{max}}/3=5\times 10^{-4}$ (wackerly2008mathematical). The resulting covariance is $\mathbf{R}_{\mathrm{sensor}}=\sigma^{2}=2.5\times 10^{-7}$. Note that we use the Joseph form (bucy2005filtering) for the covariance update in Eq.˜15, for improved stability. Preliminary analysis of the validation trajectory showed that the average ratio of $\frac{\mathbf{W}_{k}}{\mathbf{R}_{k}}$ was high, causing the filter to overtrust measurement updates and yield noisy state estimates. To address this, $\mathbf{W}_{k}$ was decreased by a factor of 100. After each update, the PINN initial state is reset to $\hat{\mathbf{x}}_{k|k}$ to ensure that the state transition model starts from the most recent state estimate. We apply the filter for each model in the ensemble separately, and we only use the ensemble mean of state estimates for the process noise covariance calculations.
@@ 5.3 Results
This section presents the results of the state estimation analysis. First, an initial guess for the phase heights is obtained by uniformly sampling 100 height vectors within their bounds and using those as initial state inputs to PINN (and VNN) to predict the outlet volume flows at $\tau=0$. These predicted outflows are compared against the experimentally measured outlet flows at $\tau=0$, and the candidate yielding the best agreement is selected as the initial state estimate. We then feed these initial states into the filter and proceed to track and estimate the phase heights for each PINN (and VNN) model individually. The individual trajectories, along with the ensemble means, are shown in Figs. 12 and 13. As in Section˜5, model performance and uncertainty are quantified using RMSE and time-averaged IQR, respectively.
In Fig.˜12, we report the results for the interpolation test trajectory. For both PINN and VNN, the estimated initial DPZ height $h_{\textrm{DP}}(\tau_{0})$ deviates from the experimental trajectory, indicating weak reconstruction of the internal states from outlet flow measurements alone. Nevertheless, in both cases the ensemble mean converges to the experimental trajectory at around $\tau=$100\text{\,}\mathrm{s}\text{/}$\$. For the heavy phase height, the PINN consistently outperforms the VNN. Despite the significant overshoot observed in the forward simulation results shown in Fig.˜10(b), the VNN–filter combination achieves comparable accuracy to the PINN ensemble in predicting DPZ height for the interpolation trajectory. This indicates that the filter update step effectively compensates for the VNN’s error accumulation by incorporating outflow measurements. However, compared to the PINN, the individual estimates of the VNN ensemble exhibit lesser consistency, as shown by the lower IQR value of the PINN ensemble.
For the extrapolation test trajectory in Fig.˜13, the PINN ensemble performs better than the VNN ensemble in case of the DPZ height. In particular, after the maximum control action (an operating value unseen in the training trajectory) is applied at around process time $\tau=$500\text{\,}\mathrm{s}\text{/}$\$, the VNN mean estimate overshoots the experimental DPZ height, whereas the PINN continues to track the DPZ height with good accuracy. Again, the individual estimates from the PINN ensemble are also more consistent than the ones from the VNN, as evidenced by the lower IQR value. For the heavy phase height, the VNN and PINN ensembles demonstrate similar accuracy, with the PINN underestimating the heavy phase height around process time $\tau=$500\text{\,}\mathrm{s}\text{/}$\$, while the VNN shows increasing deviation from the experimental trajectory thereafter.
@@
Figure 12: State estimation results for the interpolation test trajectory. Estimations from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line. The reported RMSE values quantify the deviation between the ensemble mean estimation of each model and the corresponding experimental trajectory over the entire trajectory. Time-averaged IQR values are reported to quantify the ensemble spread.
@@
Figure 13: State estimation results for the extrapolation test trajectory. Estimations from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line. The reported RMSE values quantify the deviation between the ensemble mean estimation of each model and the corresponding experimental trajectory over the entire trajectory. Time-averaged IQR values are reported to quantify the ensemble spread.
After estimating the average DPZ height along the separator using the filter, we apply the simple NN to predict the DPZ height at the separator end (Fig.˜14). Both the PINN and VNN ensembles demonstrate good accuracy. Deviations occur primarily when the average DPZ height itself deviates, indicating that the NN performs well, but errors propagate where the underlying ensemble predictions are inaccurate (cf. Figs.˜12 and 13). This effect is emphasized in the extrapolation trajectory (cf. Fig.˜13), where larger deviations in the VNN ensemble lead to correspondingly larger errors in the predicted DPZ height at the separator end.
@@
Figure 14: Prediction of the DPZ height at the separator end. Experimental results are obtained from camera QIR04 at detection window $h_{4,3}$. Estimates of the average DPZ from both PINN and VNN are fed into the NN that predicts the DPZ height at separator end. The reported RMSE values quantify the deviation between the prediction of each model and the corresponding experimental trajectory over the entire trajectory.
We therefore conclude that the two-stage trained PINN, when combined with the filter, enables effective real-time estimation of the internal separator states based on outlet flow measurements. As the outlet flow-rate measurements are inexpensive to obtain, our PINN constitutes a practical soft sensor for the difficult-to-measure phase heights. Moreover, we demonstrate that predicting the DPZ height at the separator end (the maximum value) from the average DPZ height is possible using a simple NN.
@@ 6 Conclusion and Outlook
We investigated the estimation of phase heights in a pilot-scale liquid–liquid separator operating with 1-octanol and deionized water at 30 °C under varying flow conditions. Direct measurement of phase heights is challenging, as the interior of the separator is not readily accessible for reliable observation of the phase interface. Although we employ a YOLOv8-based phase-boundary detection algorithm in our experimental setup, the resulting measurements still contain noise due to the difficulty of distinguishing the interface and the sensitivity of the algorithm to illumination conditions. Moreover, the available mechanistic model relies on multiple simplifying assumptions, resulting in limited predictive accuracy. To address these challenges, we developed a PINN-based dynamic model and proposed a two-stage training strategy: pretraining on approximate physics, i.e., synthetic data from the low-fidelity model, followed by fine-tuning with experimental data. We further used ensemble learning to mitigate the influence of outlier models. Moreover, the ensemble not only improves prediction robustness but also provides a practical measure of model uncertainty, which is incorporated into the state estimation framework through adaptive covariance estimation. Finally, we combined the trained PINN with an EKF-inspired filter to enable estimation of the aqueous phase and DPZ heights. The filter prevents error accumulation from chaining of PINN predictions by fusing model predictions with real-time flow measurements to update the estimated phase-height trajectories.
@@
The two-stage PINN ensemble outperformed both the low-fidelity mechanistic model and a single-stage PINN ensemble trained only on experimental data, showing that integrating mechanistic model knowledge with experimental measurements yields better results than either approach alone. For further comparison, we also pretrained and fine-tuned an ensemble of purely data-driven vanilla neural networks (VNNs). Both the PINN and VNN ensembles were embedded in a state estimation framework to enable phase height estimation during deployment from flow-rate measurements only. The PINN ensemble achieved higher accuracy in tracking phase heights. Furthermore, a simple neural network can predict the outlet DPZ height (typically the maximum value and most critical for flooding) from average DPZ height estimates. The results demonstrate that the PINN-based state estimation can serve as a reliable soft sensor, enabling the real-time monitoring of this separation process from readily available flow-rate measurements.
@@
A few limitations of the present study should be noted. First, the model assumes a band-shaped DPZ, which becomes inaccurate at higher flow rates where the DPZ develops a distinct wedge-shaped profile. Although a mapping between the average and maximum DPZ height was learned by a simple neural network, the PINN model does not fully represent the actual separator dynamics. Second, the experimental data contain substantial gaps caused by missing phase-boundary detections, particularly for QIR03, which limits the training data quality. Third, the experimental dataset includes no trajectories with conditions close to flooding. While such conditions are difficult to obtain experimentally for safety reasons, their absence restricts the ability of the model to detect or classify flooding behavior. Fourth, we acknowledge that filtering the immeasurable internal flows in the proposed state estimation framework could improve phase height estimation accuracy, however this is not possible with the current PINN architecture. Lastly, the experimental dataset comprises only four trajectories, with only one being used for training.
@@ Acknowledgements
During the preparation of this work, generative AI tools (ChatGPT, Claude) were used to assist with grammar, spelling, and style improvements in the writing. Generative AI coding tools (Claude) were used for post-processing tasks, specifically for plotting and visualization of the results. In all cases, the authors reviewed and edited the outputs, and they take full responsibility for the integrity and accuracy of the publication.